# Understanding Literary Texts by LLMs: A Case Study of Ancient Chinese Poetry

### v1

@@ 4.3 Probability
Another metric that captures the novelty and originality of a poetic composition is the absolute probability [^19] For ease of reading, we convert all probabilities into percentages. of its constituent tokens. In this paper, we assemble a world probability distribution  [^20] https://github.com/ByteDance-high-level-creation/literary_appreciation/blob/master/world_token_probs.npy for all tokens, drawing from the BAAI-CCI 2.0 dataset  [^21] https://data.baai.ac.cn/details/BAAI-CCI2. Additionally, we procure and integrated probability distributions specific to ancient Chinese  [^22] https://github.com/ByteDance-high-level-creation/literary_appreciation/blob/master/ancient_token_probs.npy, Qilv  [^23] https://github.com/ByteDance-high-level-creation/literary_appreciation/blob/master/qilv_token_probs.npy, and Ci  [^24] https://github.com/ByteDance-high-level-creation/literary_appreciation/blob/master/ci_token_probs.npy. The synthesis of these datasets has enabled us to derive an absolute probability table that is pertinent to the poetic genre.

### v2

@@ 4.3 Probability
Another metric that captures the novelty and originality of a poetic composition is the absolute probability [^19] For ease of reading, we convert all probabilities into percentages. of its constituent tokens. In this paper, we assemble a world probability distribution  [^20] https://github.com/ByteDance-high-level-creation/literary_appreciation/blob/master/world_token_probs.json for all tokens, drawing from the BAAI-CCI 2.0 dataset  [^21] https://data.baai.ac.cn/details/BAAI-CCI2. Additionally, we procure and integrated probability distributions specific to ancient Chinese  [^22] https://github.com/ByteDance-high-level-creation/literary_appreciation/blob/master/ancient_token_probs.json, Qilv  [^23] https://github.com/ByteDance-high-level-creation/literary_appreciation/blob/master/qilv_token_probs.json, and Ci  [^24] https://github.com/ByteDance-high-level-creation/literary_appreciation/blob/master/ci_token_probs.json. The synthesis of these datasets has enabled us to derive an absolute probability table that is pertinent to the poetic genre.
@@
Table 12: The KL divergence between output probability of the base model and absolute probability for different authors in the Tang and Song dynasties
@@
Table 13: The KL divergence between output probability of the base model and absolute probability for different authors in the Ming and Qing dynasties